1. Find the projection and reflection of the point (1,2,3) with respect to the plane defined by x-3y+z=30
Answer #1

Solution-

. Let the equation of a line passes through the point A(1,2,3) and perpendicular to the plane n-39+2 = 30 is n-I =y=2=2-3 = L